OpenStudy (anonymous):

multiply: (1+2i)(2-i)

OpenStudy (anonymous):

i am so lost on where to start with this problem

OpenStudy (anonymous):

this is 8th grade math right?

OpenStudy (anonymous):

intermediate algebra. would i answer it the same if it was 'x' instead of 'i'

OpenStudy (watchmath):

You just apply the distributive property \(1\cdot 2+1\cdot (-i)+2i\cdot2+2i\cdot (-i)\)

OpenStudy (anonymous):

Just foil it out like you would normally (treat the i kinda like a variable) and just remember that if you have an \(i^2\) that equals -1

OpenStudy (anonymous):

oh okay!! thanks! i wasnt sure if i just fouil like normal. the "i" threw me off!

OpenStudy (anonymous):

don't be scared of letters. They're actually much easier to work with than numbers because you don't have to plug them into your calculator ;)

OpenStudy (anonymous):

thank you. i got 2-1i+4i-2i so far

OpenStudy (anonymous):

2+3i-2i as the answer

OpenStudy (watchmath):

\(2-1i+4i-2i^2\). Now you can use \(i^2=-1\)

OpenStudy (anonymous):

2+3i-2i^2 ..which would be 2+3i+2

OpenStudy (watchmath):

good! you can combine 2+2 as well